### v1

@@ Abstract
which is shown in Fig. 3(a). Similarly, we construct the Wiener filter for $\hat{p}$, and compute the conditional momentum spectrum $S_{\delta p\delta p}$ and cross spectrum $S_{\delta q\delta p}$ (see Methods E). As we can see in Fig. 3(a), the conditional spectrum is strongly suppressed compared to the measured and modeled spectra. To account for the diverging displacement energy at low frequency given by the structural damping model (see Methods F and [33]), we calculate the covariances by integrating the conditional (cross) spectra from $10$ kHz to $1$ MHz to avoid the divergence, e.g. the cross correlation between $\delta\hat{q}$ and $\delta\hat{p}$ is $C_{\delta q\delta p}=\int_{2\pi\times 10^{4}\,\mathrm{Hz}}^{2\pi\times 10^{6}\,\mathrm{Hz}}\frac{d\Omega}{2\pi}\mathrm{Re}[S_{\delta q\delta p}[\Omega]]$. Normalizing the (co)-variances to quanta in displacement $2q_{\mathrm{zpf}}^{2}=\hbar/m\Omega_{\mathrm{eff}}$ and momentum $2p_{\mathrm{zpf}}^{2}=\hbar m\Omega_{\mathrm{eff}}$, the covariance matrix of the estimation error is
$\tilde{\mathbb{V}}=\begin{pmatrix}\tilde{V}_{\delta q\delta q}&\tilde{C}_{\delta q\delta p}\\ \tilde{C}_{\delta q\delta p}&\tilde{V}_{\delta p\delta p}\end{pmatrix}=\begin{pmatrix}0.88&1.04\\ 1.04&3.61\end{pmatrix},$ (7)
where the tilde represents normalization to the energy quanta. The minimum eigenvalue is $0.53$, only $(0.28\pm 0.18)$ dB above ($1.07\pm 0.04$ times) the ZPF ($0.5$), and the maximum eigenvalue is $3.96$, $(8.98\pm 0.11)$ dB above ($7.9\pm 0.2$ times) the ZPF level. To properly evaluate the influence of uncertainties of experimental parameters on the conditional variances, the mean values and uncertainties reported here come from Monte Carlo uncertainty propagation. As a property of Gaussian states, the conditional covariance matrix $\mathbb{V}$ and the mean values of the estimators $\langle\hat{q}_{\mathrm{e}}\rangle$ and $\langle\hat{p}_{\mathrm{e}}\rangle$ define the conditional state uniquely, and the uncertainty of the state is given by the covariance matrix. The quadrature uncertainties corresponding to $\tilde{V}$ are shown in Fig. 3(c), with 1 standard deviation uncertainty displayed. Our result is the highest quality realization of the protocol to date, reducing the smallest variance by 3 orders of magnitude compared to the previous record [32].
Figure 3: a) Amplitude spectral densities of the directly measured homodyne photocurrent, the calibrated modeled homodyne photocurrent, and the displacement conditional spectrum as in Eqn. 6. b) The corresponding displacement Wiener filter as in Eqn. 5. c) Quadrature diagram of the derived conditional state. The light purple band corresponds to the region of one sigma confidence.
@@
We have demonstrated preparation of a classically squeezed mechanical state approaching the quantum regime in post processing. We cleaned the measured spectrum by comparing the target spectrum to spectra taken at different cavity detunings. The same set of data compared to different auxiliary spectra shows the validity of the theoretical model. Though data cleaning is in post processing, this procedure can be converted to a real-time filter defined by the measured spectra, similar to causal Wiener filters. Together with the causal Wiener filters, real-time state preparation will be possible.
There is, however, non-stationary noise in the measurement which masks the physical motion of the mechanical oscillator and cannot be eliminated completely by data cleaning. This non-stationary noise appears in measurements at multiple positions in the experiment, and is suspected to have electronic origin. The omnipresent feature makes it challenging to hunt down and eliminate the noise. To recover the physical displacement, we used the model to compute the causal Wiener filter and the conditional covariance matrix. This compromise prevents us from filtering the photo current to get real-time estimation of the mechanical state. Eliminating the non-stationary noises is a necessary step for quantum state preparation in real time.
Our result stands on the border line of the quantum regime. For preparation of a definitive quantum squeezed state, higher readout efficiency and lower cavity detuning (hence a lower optical spring frequency) would be helpful. The former requires optimizing the detection chain by replacing current optics with lower loss versions. The latter requires locking the optomechanical cavity on resonance, for example by using a Pound-Drever-Hall lock. Additionally, injecting light from the macroscopic mirror side of the cavity and conducting the measurement in transmission is also beneficial, as it avoids excess shot noise from the promptly reflected light that is not spatially matched to the cavity mode. With these improvements, quantum squeezed mechanical state preparation of a 50 ng mirror at room temperature should be possible [42].
Very recently, levitated particles have been prepared into quantum squeezed state in the optical regime [19]. Compared to them, our system is 7 orders of magnitude more massive. This drastic difference in mass allows our optomechanical system to explore a different region of parameter space of the classical-quantum transition, as well as to probe gravitational effects on quantized motion [10]. The steady state controls used in the continuous measurement make it straightforward to adapt these protocols to the LIGO test masses. With feedback stiffening [41] and improvements in sensitivity around $100$ Hz [43], preparing the 10-kg-scale LIGO test masses in a quantum squeezed state should be possible in near future. This will be a great leap forward in studying the problem of quantum-to-classical transition. LIGO test masses will also become the most sensitive quantum enhanced accelerometers ever, which will contribute to dark matter searches [44].
@@ Acknowledgements
  • Data and code availability: Data and code underlying the results presented in this paper are not publicly available at this time but may be obtained from the authors upon request.
  • Author contribution: B.B.L. and J.C. constructed the theoretical model, conducted the experiment and analyzed the data; R.E.P., S.A., X.Y. contributed to the preparation of the experiment and data taking; G.D.C. fabricated the cantilever chip; X.Y., T.R.C. and N.M. supervised the project.
@@
where $G=\partial\Omega_{\mathrm{cav}}/\partial\hat{q}$ is the optomechanical interaction strength, with $\Omega_{\mathrm{cav}}$ the cavity angular resonance frequency; $\alpha=\sqrt{n_{\mathrm{cav}}}$ is the classical amplitude of the intra-cavity optical field, with $n_{\mathrm{cav}}$ the intra-cavity photon number; $\kappa_{1}$ is the angular cavity coupling rate on the cantilever side, $\kappa_{2}$ the angular cavity loss rate on the curved mirror side, and $\kappa_{\mathrm{L}}$ the intra-cavity angular loss rate; $\hat{X}_{\mathrm{in}1}$, $\hat{X}_{\mathrm{in}2}$, and $\hat{X}_{\mathrm{inL}}$ are the input amplitude fluctuations from the cantilever, curved mirror and internal loss; The factors $u[\Omega]=(\kappa/2+i\Omega)/((\kappa/2+i\Omega)^{2}+\Delta^{2})$, $v[\Omega]=-\Delta/((\kappa/2+i\Omega)^{2}+\Delta^{2})$ are the components of the cavity susceptibility, where $\kappa$ is the total angular linewidth of the cavity, and $\Delta$ the angular cavity detuning; $\hat{\xi}_{\mathrm{th}}$ is the thermal force operator. $\chi_{\mathrm{m}}[\Omega]=1/(m(\Omega_{\mathrm{m}}^{2}-\Omega^{2}+i\Gamma_{\mathrm{m}}[\Omega]\Omega))$ is the bare mechanical susceptibility, with $m$, $\Omega_{\mathrm{m}}$, and $\Gamma_{\mathrm{m}}[\Omega]$ the effective mass, the natural angular frequency of the mechanical mode, and the natural mechanical linewidth, respectively. Here the frequency dependence of $\Gamma_{\mathrm{m}}$ is a result of structural damping, which will be discussed in detail in Methods F. In most literature, the optomechanical interaction strength is rewritten in terms of $g=G\alpha x_{\mathrm{zpf}}$, with $x_{\mathrm{zpf}}=\sqrt{\hbar/2m\Omega_{\mathrm{eff}}}$ the ZPF of displacement, where $\Omega_{\mathrm{eff}}$ is the effective mechanical angular frequency influenced by optical spring and feedback. However, we intentionally keep the above form without ZPF, as $\Omega_{\mathrm{eff}}$ is strongly influenced by optical springing in our system. Keeping $G$ explicitly gives us an expression that works for arbitrary optical spring [45].
@@
where the last term is given by feedback. Collecting $\hat{X}[\Omega]$ on both sides, we get the close-loop expression of $\hat{X}[\Omega]$:
@@
Similarly, the close-loop $\hat{Y}[\Omega]$ is
@@
where $\hat{z}$ is the imprecision noise in displacement unit. $\hat{y}$ is sometimes referred as the apparent displacement.
Methods B Optical Parameter Characterization
@@
We use the optical spring to constrain the optical parameters, as in [36]. The optical spring frequency is determined by measuring the open-loop cavity transfer function from amplitude fluctuations to photo current. The measurement scheme is shown in Fig. 5. A Newport LB1005 servo is used to lock the cavity, and as a sum point of signal and drive. A Zürich Instrument HF2 lock-in amplifier is used as the vector network analyzer. The signals right before and after the servo are fed to the two inputs A and B of the HF2, where the ratio of the two is computed. The transfer function of the amplitude modulator is flat in the frequency range we are interested in. Therefore, it only contributes a constant factor to the loop, which is absorbed to the controller $K$ for simplicity. The close-loop inputs to Port A and B of the HF2 are given by
@@
Methods D Noisy Background Cancellation
The direct spectra of the homodyne photo currents contain noises not included in the theoretical model, such as classical laser phase noise, higher order mechanical modes of the cantilever, mechanical motion of the structure holding the optomechanical cavity, noise from the piezo used to lock the cavity, calibration tones etc. Here we introduce a noise cancellation procedure to recover the spectrum of the fundamental mechanical mode and quantum shot noise from the noisy environment. The basic idea is to use homodyne spectra measured with different experimental parameters, especially different cavity detunings, to eliminate the stationary excess noises in the target spectrum. In the following, we refer the spectra used to clean the target spectrum as auxiliary spectra.
@@
Figure 8 illustrates our noise cancellation procedure. We took homodyne photo currents for different cavity detunings, the sample spectra of which are shown in Fig. 8(a). The sharp peaks at low frequencies aligning with integer number of 10 kHz are the piezo calibration tone at 10 kHz and its harmonics. At 60 kHz a phase calibration tone overlaps with harmonics of the piezo calibration tone. At non-integer 10 kHz frequencies, the sharp peaks are high order mechanical modes, which couple weakly to the cavity field, and do not experience significant optical springing. The broad structure around 28 kHz corresponds to the cavity piezo resonance. The common sharp structures above 130 kHz are phase noises from the laser and fibers. Figure 8(b) shows the spectra of the lowest detuning measurement cleaned by 4 different auxiliary spectra. All spectra match the modelled curve well. We notice that around the optical springed mechanical resonance, there are some noise peaks that are not consistent among the spectra. These structures correspond to non-stationary electronic noise. The harmonics of the piezo calibration tone are mitigated in the cleaned spectra, but not eliminated, as the nonlinearity of the piezo response depends on the DC bias, which can differ from measurement to measurement. Due to these non-stationary noises, we calculate the final mechanical squeezing using the model instead of the cleaned data.
Methods E Momentum and cross spectra of the fundamental mechanical mode
@@
Methods F Structural Damping and Choice of Integration Band
@@
$\chi_{\mathrm{m}}[\Omega]=\frac{1}{m(\Omega_{\mathrm{m}}^{2}-\Omega^{2}+i\Omega_{\mathrm{m}}^{2}\phi_{\mathrm{l}})}=\frac{1}{m(\Omega_{\mathrm{m}}^{2}-\Omega^{2}+i\Gamma_{\mathrm{m}}[\Omega]\Omega},$ (35)

### v2

@@ Abstract
which is shown in Fig. 3(a). Similarly, we construct the Wiener filter for $\hat{p}$, and compute the conditional momentum spectrum $S_{\delta p\delta p}$ and cross spectrum $S_{\delta q\delta p}$ (see Methods F). As can be seen in Fig. 3(a), the conditional spectrum is strongly suppressed compared to the measured and modeled spectra. To account for the diverging displacement energy at low frequency given by the structural damping model (see Methods G and [34]), we calculate the covariances by integrating the conditional (cross) spectra from $10$ kHz to $1$ MHz to avoid the divergence, e.g. the cross correlation between $\delta\hat{q}$ and $\delta\hat{p}$ is $C_{\delta q\delta p}=\int_{2\pi\times 10^{4}\,\mathrm{Hz}}^{2\pi\times 10^{6}\,\mathrm{Hz}}\frac{d\Omega}{2\pi}\mathrm{Re}[S_{\delta q\delta p}[\Omega]]$. Normalizing the (co)-variances to ZPF in displacement $q_{\mathrm{zpf}}^{2}=\hbar/2m\Omega_{\mathrm{eff}}$ and momentum $p_{\mathrm{zpf}}^{2}=\hbar m\Omega_{\mathrm{eff}}/2$, the covariance matrix of the estimation error is
$\tilde{\mathbb{V}}=\begin{pmatrix}\tilde{V}_{\delta q\delta q}&\tilde{C}_{\delta q\delta p}\\ \tilde{C}_{\delta q\delta p}&\tilde{V}_{\delta p\delta p}\end{pmatrix}=\begin{pmatrix}1.76&2.08\\ 2.08&7.22\end{pmatrix},$ (7)
where the tilde represents normalization to the ZPF. The minimum eigenvalue is $1.07\pm 0.04$, only $(0.28\pm 0.18)$ dB above the ZPF, and the maximum eigenvalue is $7.9\pm 0.2$, $(8.98\pm 0.11)$ dB above the ZPF level. To accurately assess the impact of experimental parameter uncertainties on the conditional variances, the reported mean values and uncertainties are obtained through Monte Carlo uncertainty propagation (see Methods H). As a property of Gaussian states, the conditional covariance matrix $\mathbb{V}$, together with the mean values of the estimators $\langle\hat{q}_{\mathrm{e}}\rangle$ and $\langle\hat{p}_{\mathrm{e}}\rangle$, uniquely defines the conditional state. The state’s uncertainty is fully characterized by the covariance matrix. The quadrature uncertainties corresponding to $\tilde{V}$ are shown in Fig. 3(c), with 1 standard deviation uncertainty displayed for reference. Our result is the highest quality realization of the protocol to date, reducing the smallest variance by three orders of magnitude compared to the previous record [33].
Figure 3: a) Amplitude spectral densities of the directly measured homodyne photocurrent, the calibrated modeled homodyne photocurrent, and the conditional displacement spectrum as in Eqn. 6. b) The corresponding displacement Wiener filter as defined in Eqn. 5. c) Quadrature diagram of the derived conditional state. The light purple band indicates the one-sigma confidence region.
@@
We have demonstrated the preparation of a classically squeezed mechanical state approaching the quantum regime using a measurement-based protocol. The measured spectrum was cleaned by comparing the target spectrum to auxiliary spectra acquired at different cavity detunings. Applying this procedure to the same dataset with various auxiliary spectra confirms the validity of the theoretical model. Although the data cleaning was performed after the measurement was recorded, this approach can be implemented as a real-time filter based on the measured spectra, analogous to causal Wiener filters. Combined with causal Wiener filtering, this enables the prospect of real-time state preparation.
However, our current setup still falls short of preparing a quantum squeezed mechanical motional state in real-time. There are some key further steps required to achieve the goal. First, we need to eliminate non-stationary noise, which prevents direct filtering of the photocurrent for realtime state estimation. Second, a definitive quantum state could be achieved if we could operate the optical cavity closer to resonance, a step that requires a major reworking of the control systems in our setup. Third, we need to improve the readout efficiency, which would be achieved by reducing optical losses throughout the experiment and also increasing the photodetection efficiency. Each of these represents a significant technical upgrade in our current setup, and are the subject of future work, with more details provided in Methods I. With these improvements, the preparation of a quantum-squeezed mechanical state of our 50 ng mirror at room temperature should be achievable [43].
Recently, levitated particles have been prepared in quantum-squeezed states in the optical regime [20, 21]. Compared to these systems, our optomechanical platform is seven orders of magnitude more massive. This substantial difference in mass enables exploration of a distinct region of the parameter space associated with the quantum-to-classical transition, as well as the potential to probe gravitational effects on quantized motion [7].
Even when operating in thermal states and with non-quantum-limited readout, cantilever systems remain among the most promising platforms for investigating the quantum-to-classical transition, provided that suitable materials and structural designs are employed [44, 45]. Our work advances this effort by enabling significantly more precise tests using potential quantum states and quantum-limited readout. The steady-state control strategies used in continuous measurement also make it straightforward to adapt these protocols to LIGO test masses.
With feedback stiffening [42] and ongoing improvements in sensitivity around $100$ Hz [46], preparing the 10-kg-scale LIGO test masses in a quantum-squeezed state should be feasible in the near future. Such an achievement would represent a major milestone in the study of the quantum-to-classical transition. Furthermore, LIGO test masses would become the most sensitive quantum-enhanced accelerometers ever developed, directly contributing to dark matter searches [47].
@@ Acknowledgements
  • Data and code availability: Source data for Fig. 3 is available at https://doi.org/10.7910/DVN/9FGAQF. Raw data and code underlying the results presented in this paper may be obtained from the authors upon request.
  • Author contribution: B.B.L. and J.C. constructed the theoretical model, conducted the experiment and analyzed the data; R.E.P., S.A., X.Y. contributed to the preparation of the experiment and data taking; G.D.C. designed and fabricated the cantilever chip; X.Y., T.R.C. and N.M. supervised the project.
@@
where $G=\partial\Omega_{\mathrm{cav}}/\partial\hat{q}$ is the optomechanical interaction strength, with $\Omega_{\mathrm{cav}}$ the cavity angular resonance frequency; $\alpha=\sqrt{n_{\mathrm{cav}}}$ is the classical amplitude of the intra-cavity optical field, with $n_{\mathrm{cav}}$ the intra-cavity photon number; $\kappa_{1}$ is the angular cavity coupling rate on the cantilever side, $\kappa_{2}$ the angular cavity loss rate on the curved mirror side, and $\kappa_{\mathrm{L}}$ the intra-cavity angular loss rate; $\hat{X}_{\mathrm{in}1}$, $\hat{X}_{\mathrm{in}2}$, and $\hat{X}_{\mathrm{inL}}$ are the input amplitude fluctuations from the cantilever, curved mirror and internal loss; the factors $u[\Omega]=(\kappa/2+i\Omega)/((\kappa/2+i\Omega)^{2}+\Delta^{2})$, $v[\Omega]=-\Delta/((\kappa/2+i\Omega)^{2}+\Delta^{2})$ are the components of the cavity susceptibility, where $\kappa$ is the total angular linewidth of the cavity, and $\Delta$ the angular cavity detuning; $\hat{\xi}_{\mathrm{th}}$ is the thermal force operator. $\chi_{\mathrm{m}}[\Omega]=1/(m(\Omega_{\mathrm{m}}^{2}-\Omega^{2}+i\Gamma_{\mathrm{m}}[\Omega]\Omega))$ is the bare mechanical susceptibility, with $m$, $\Omega_{\mathrm{m}}$, and $\Gamma_{\mathrm{m}}[\Omega]$ the effective mass, the natural angular frequency of the mechanical mode, and the natural mechanical linewidth, respectively. Here the frequency dependence of $\Gamma_{\mathrm{m}}$ is a result of structural damping, which will be discussed in detail in Methods G. In most literature, the optomechanical interaction strength is rewritten in terms of $g=G\alpha x_{\mathrm{zpf}}$, with $x_{\mathrm{zpf}}=\sqrt{\hbar/2m\Omega_{\mathrm{eff}}}$ the ZPF of displacement, where $\Omega_{\mathrm{eff}}$ is the effective mechanical angular frequency influenced by optical spring and feedback. However, we intentionally keep the above form without ZPF, as $\Omega_{\mathrm{eff}}$ is strongly influenced by optical springing in our system. Keeping $G$ explicitly gives us an expression that works for arbitrary optical spring [48].
@@
where the last term is given by feedback. Collecting $\hat{X}[\Omega]$ on both sides, we get the closed-loop expression of $\hat{X}[\Omega]$:
@@
Similarly, the closed-loop $\hat{Y}[\Omega]$ is
@@
where $\hat{z}$ is the imprecision noise in displacement unit. $\hat{y}$ is sometimes referred to as the apparent displacement.
Methods B State Estimation and Conditional State
For Gaussian states, the mean values of the estimators $\langle\hat{q}_{\mathrm{e}}\rangle$ and $\langle\hat{p}_{\mathrm{e}}\rangle$ and the (co)-variances of their corresponding errors $V_{\delta q\delta q}$, $V_{\delta p\delta p}$ and $C_{\delta q\delta p}$ uniquely define a state centered at the mean values in phase space with uncertainties given by the estimation errors [26, 27, 31]. As the estimators and estimation errors depend on the measurement record $\hat{y}$, the state is deemed a conditional state. The mean values directly inherit the stochastic nature of the measurement record $\hat{y}(t)$, resulting in a random walk of the center of the conditional state in phase space, within the distribution of a thermal state given by the thermal and back-action baths, as shown in Fig. 1(a) in the main text. In contrast, the variances of the estimation errors decrease deterministically in time and converge to a steady-state value, as more information is extracted from the measurement process [30]. The steady-state value is determined by the measurement rate $\Gamma_{\mathrm{meas}}$, the thermal decoherence rate $\Gamma_{\mathrm{th}}$ and the back-action rate $\Gamma_{\mathrm{ba}}$ shown in Fig. 1(b) and the mechanical frequency $\Omega_{\mathrm{m}}$ shown in Fig. 1(c). A conditional state prepared by causal filters (constructed using past information) can be transferred to an unconditional state through feedback in real time [32]. On the other hand, due to the requirement of future information, a conditional state prepared by non-causal filters is not accessible in real time, and cannot be transferred to an unconditional state by feedback. Moreover, non-causal filters may result in unphysical conditional states, which violates the Heisenberg uncertainty principle [49]. Therefore, for measurement-based preparation of physical mechanical states, estimate filters should be causal. In addition, to minimize the uncertainty, causal Wiener filters are the typical choices of $H_{q}$ and $H_{p}$. We follow this practice in our work.
Methods C Optical Parameter Characterization
@@
We use the optical spring to constrain the optical parameters, as in [37]. The optical spring frequency is determined by measuring the open-loop cavity transfer function from amplitude fluctuations to photocurrent. The measurement scheme is shown in Fig. 5. A Newport LB1005 servo is used to lock the cavity, and as a sum point of signal and drive. A Zürich Instrument HF2 lock-in amplifier is used as the vector network analyzer. The signals right before and after the servo are fed to the two inputs A and B of the HF2, where the ratio of the two is computed. The transfer function of the amplitude modulator is flat in the frequency range we are interested in. Therefore, it only contributes a constant factor to the loop, which is absorbed to the controller $K$ for simplicity. The closed-loop inputs to Port A and B of the HF2 are given by
@@
Methods E Noisy Background Removal
The direct spectra of the homodyne photocurrents contain noises not included in the theoretical model, such as classical laser phase noise, higher order mechanical modes of the cantilever, mechanical motion of the structure holding the optomechanical cavity, noise from the piezo used to lock the cavity, calibration tones etc. Here we introduce a noise removing procedure to recover the spectrum of the fundamental mechanical mode and quantum shot noise from the noisy environment. The basic idea is to use homodyne spectra measured with different experimental parameters, especially different cavity detunings, to eliminate the stationary excess noises in the target spectrum. In the following, we refer the spectra used to clean the target spectrum as auxiliary spectra.
@@
Figure 8 illustrates our noise removal procedure. We took homodyne photocurrents for different cavity detunings, the sample spectra of which are shown in Fig. 8(a). The sharp peaks at low frequencies aligning with integer number of 10 kHz are the piezo calibration tone at 10 kHz and its harmonics. At 60 kHz a phase calibration tone overlaps with harmonics of the piezo calibration tone. At non-integer 10 kHz frequencies, the sharp peaks are high order mechanical modes, which couple weakly to the cavity field, and do not experience significant optical springing. The broad structure around 28 kHz corresponds to the cavity piezo resonance. The common sharp structures above 130 kHz are phase noises from the laser and fibers. Figure 8(b) shows the spectra of the lowest detuning measurement cleaned by 4 different auxiliary spectra. All spectra match the modelled curve well. We notice that around the optical springed mechanical resonance, there are some noise peaks that are not consistent among the spectra. These structures correspond to non-stationary electronic noise. The harmonics of the piezo calibration tone are mitigated in the cleaned spectra, but not eliminated, as the nonlinearity of the piezo response depends on the DC bias, which can differ from measurement to measurement. Due to these non-stationary noises, we calculate the final mechanical squeezing using the model instead of the cleaned data.
Methods F Momentum and cross spectra of the fundamental mechanical mode
@@
Methods G Structural Damping and Choice of Integration Band
@@
$\chi_{\mathrm{m}}[\Omega]=\frac{1}{m(\Omega_{\mathrm{m}}^{2}-\Omega^{2}+i\Omega_{\mathrm{m}}^{2}\phi_{\mathrm{l}})}=\frac{1}{m(\Omega_{\mathrm{m}}^{2}-\Omega^{2}+i\Gamma_{\mathrm{m}}[\Omega]\Omega)},$ (35)
@@
Methods H Monte Carlo Uncertainty Propagation
Our causal Wiener filter and conditional state are calculated using the theoretical model which is based on experimental parameters. The uncertainties in these parameters affect the final conditional state. Therefore, it is crucial to study the impact of these uncertainties. Given the complicated model and filtering process, Monte Carlo uncertainty propagation is the most straightforward way to quantify the impact.
The uncertainties of optical parameters such as $\kappa$, $\delta_{\mathrm{cav}}$, $P_{\mathrm{cav}}$ come from the least-square fitting algorithm introduced in Method C. The uncertainty of the temperature is estimated from the temperature range displayed by the thermometer during the experiment. We assume that the uncertainties follow Gaussian distributions and pass the mean values and uncertainties of the parameters to the Monte Carlo algorithm, where 985 samples were randomly drawn. Due to the non-linearity in the propagation, the means of the covariances of these samples slightly deviates from the values calculated from the parameter means directly. We choose to report the Monte Carlo mean. The choice does not change the qualitative conclusion of our work.
Methods I Future Improvements to Achieve Real-Time Quantum State Preparation
Our current setup falls short of preparing a quantum squeezed mechanical motional state in real-time. The final steps required to achieve the goal are outlined below.
  1. 1.
    Elimination of Non-Stationary Noise:
    Non-stationary noise in the measurement partially obscures the physical motion of the mechanical oscillator and cannot be fully removed through data cleaning. This noise is observed at multiple locations within the experimental setup and is suspected to originate from electronic sources. Its pervasive nature makes it challenging to identify and eliminate. Currently, recovery of the physical displacement relies on theoretical modeling to compute the causal Wiener filter and the conditional covariance matrix. However, this approach prevents direct filtering of the photocurrent for real-time state estimation. Eliminating non-stationary noise is therefore a critical prerequisite for real-time quantum state preparation.
  2. 2.
    Reducing Cavity Detuning:
    Our results lie at the boundary of the quantum regime. Achieving a definitive quantum-squeezed state can be achieved by reducing the optical spring frequency, which can be realized by locking the optomechanical cavity closer to the resonance, for example using a Pound–Drever–Hall (PDH) locking scheme. The challenges of applying PDH locking scheme to this system remains unknown due to the self-locking effect [41].
  3. 3.
    Enhancing Readout Efficiency :
    Reading out efficiency is another limiting factor of our results, improving which is also a path towards definitive quantum squeezing. A common method is
    • Optimizing the Optical Injection and Detection Configuration: Injecting light from the macroscopic mirror side of the cavity and performing the measurement in transmission can mitigate excess shot noise arising from promptly reflected light that is not spatially matched to the cavity mode. This effectively improves the readout efficiency and helps optimize homodyne visibility.
    • Low-loss optics: readout efficiency can be further improved by optimizing the detection chain and replacing current optical components with lower-loss alternatives.
With these improvements, the preparation of a quantum-squeezed mechanical state of a 50 ng mirror at room temperature should be achievable [43].